OpenStudy (adamaero):

Just how to start integral of ln(1+x^3)? Medal~always

OpenStudy (adamaero):

@waterineyes @johnweldon1993 @Jhannybean

OpenStudy (johnweldon1993):

Looks gross lol, I would do it by parts but I can see it'll probably get lengthy u = ln(1 + x^3) dv = dx lets see if Jhanny has any better ideas :)

OpenStudy (jhannybean):

\[\int \ln(1+x^3)dx\] Yeah.. I would try it by parts too.

OpenStudy (jhannybean):

Wow I am trying to solve it and it's extremely long -.-

OpenStudy (jhannybean):

@johnweldon1993 is right, working on it by parts works the best.

OpenStudy (johnweldon1993):

Eww this is terrible lol...eventually looks like it might need a trig sub too

OpenStudy (adamaero):

It says to approximate (from 0 to 0.5 for bounds) but I'll just ask my prof tomorrow

OpenStudy (adamaero):

thanks both

OpenStudy (jhannybean):

u = ln (1+x^3) dv = dx du = \(\frac{\sf 3x^2dx}{\sf 1+x^3}\) v = x

OpenStudy (jhannybean):

\[\int_0^{1/2} \ln(1+x^3)dx = x\ln(1+x^3) -\int_0^{1/2}x\cdot \frac{3x^2}{1+x^3}dx\]

OpenStudy (jhannybean):

For the integral part you get : \[\int \frac{x^3}{x^3+1}\] But I don't remember how to simplify this... I recall you can split it up by long division or something like that, but I forgot how to do long division -.-

OpenStudy (jhannybean):

oop, I left out the 3 after factoring it out of the integral, typo.

OpenStudy (jhannybean):

Ohh, you might have to expand it in order to solve for it. haha. Duh?

OpenStudy (fibonaccichick666):

oh crap nvm, i read it wrong and said the wrong thing

OpenStudy (fibonaccichick666):

I'm gonna delete all of it cause it is incorrect info

OpenStudy (fibonaccichick666):

so let's see if we let u=1+x^3 du=3x^2 yes?

OpenStudy (fibonaccichick666):

so then we have \[\int (3x^2)^-1 lnu du\}

OpenStudy (fibonaccichick666):

\[\int (3x^2)^{-1} ~ln~u ~~du\]

OpenStudy (fibonaccichick666):

now it is a candidate for integration by parts

OpenStudy (jhannybean):

LIPETs. log > Inverse > polys> expos> trig

OpenStudy (fibonaccichick666):

? what's that for?

OpenStudy (jhannybean):

u = ln(u) du = 1/u du dv = (3x^2)\(^{-1}\) v = ...

OpenStudy (jhannybean):

LIPETs is what I use to figure out which function I place as u and which one i place as dv

OpenStudy (fibonaccichick666):

oh, never seen it before, nice

OpenStudy (jhannybean):

But I think this case might me more complicated, so order will be switched.

OpenStudy (fibonaccichick666):

hmm, I would use that

OpenStudy (fibonaccichick666):

because integrating lnu, is not something I'd want to do

OpenStudy (jhannybean):

hmm... integrating dv.

OpenStudy (fibonaccichick666):

it's messy

OpenStudy (fibonaccichick666):

because I didn't put it in terms of u... nitwit

OpenStudy (fibonaccichick666):

u=1+x^3 right, means x=(u-1)^1/3

OpenStudy (fibonaccichick666):

should make it better

OpenStudy (fibonaccichick666):

but @adamaero, why don't you give it a shot

OpenStudy (adamaero):

no, way. Thanks all, but I've got other studying to do for everything else, and haven't started our current chapter yet, so on to that